If a car moves round a circular road of radius r at a constant speed v, (A) its velocity changes and the acceleration is v/r^2 (B) there is no force on the car since its speed is constant (C) there is no velocity change since the speed is constant (D) the force on the car is towards the center and is mv^2/r (E) the force on the car is outwards from the center and is v^2/r.
1
Expert's answer
2020-11-24T05:52:16-0500

D is the correct option.


Option A is wrong because the centripetal acceleration is "\frac{v²}r" and not "\frac{v}{r²}"


Option B is wrong because there's a centripetal force on the car since there's centripetal acceleration.


Option C is wrong because the velocity changes as the car changes direction round the circular road.


Option E is wrong because the force on the car is "m\frac{v²}r" not "\frac{v²}r"